By walking 3/4th of his usual speed, a man reaches his office 20 minutes later than usual. What is his usual time?
Answer by khwang(438)

Let L be the distance, V be his usual speed.
We have L/(3/4 V) = L/V + 20.
So, 4/3 * L/V = L/V + 20,
Or 1/3 * L/V = 20,
we get L/V = 60 (min)

Since V is his usual speed, so T = L/V his usual time for
him walking from to the office.
Another way , let T (min) be the usual time, then he spent 4 T/3
when walking in 3/4 of usual speed.
Since 4T/3 = T +20, we obtain T = 60 min.
Kenny